OpenStudy (anonymous):

A sequence is defined by the recursive formula f (n + 1) = f(n) – 2. If f(1) = 18, what is f(5)?

OpenStudy (anonymous):

f (n + 1) = f(n) – 2 means that you keep subtracting 2 to get the next number in the sequence. f(1) = 18 f(2) = 18 - 2 = 16 and so on

OpenStudy (anonymous):

I got 13

OpenStudy (anonymous):

I got 13 as well but the actual answer is 10 @thanksten

OpenStudy (anonymous):

how did you guys get 13? f(1) = 18 f(2) = 18 - 2 = 16 f(3) = 16 - 2 = 14 f(4) = 14 - 2 = 12 f(5) = 12 - 2 = 10

OpenStudy (anonymous):

Ohhhh,sorry I got 10 too

OpenStudy (anonymous):

Thanks@peachpi